In math club there are 7 girls and 5 boys. A team of 6 students must be formed to coMpete In math contest.let x be the number of boys in the team.

1. What are the posible value of x?

2. Construct a probability distribution for the random variable.

1
Expert's answer
2022-03-08T03:12:06-0500

Let's denote B - boys, G - girls.

Sample space S is all possible outcomes.

Let "X" be the random variables representing the number of number of boys in the team.

1. The possible values of "x" are "0,1,2,3,4,5."


2. There are "\dbinom{7+5}{6}=924" possible outcomes.


"P(X=0)=\dfrac{\dbinom{5}{0}\dbinom{7}{6}}{\dbinom{12}{6}}=\dfrac{1(7)}{924}=\dfrac{1}{132}"

"P(X=1)=\dfrac{\dbinom{5}{1}\dbinom{7}{5}}{\dbinom{12}{6}}=\dfrac{5(21)}{924}=\dfrac{5}{44}"

"P(X=2)=\dfrac{\dbinom{5}{2}\dbinom{7}{4}}{\dbinom{12}{6}}=\dfrac{10(35)}{924}=\dfrac{25}{66}"

"P(X=3)=\dfrac{\dbinom{5}{3}\dbinom{7}{3}}{\dbinom{12}{6}}=\dfrac{10(35)}{924}=\dfrac{25}{66}"

"P(X=4)=\dfrac{\dbinom{5}{4}\dbinom{7}{2}}{\dbinom{12}{6}}=\dfrac{5(21)}{924}=\dfrac{5}{44}"

"P(X=5)=\dfrac{\dbinom{5}{5}\dbinom{7}{1}}{\dbinom{12}{6}}=\dfrac{1(7)}{924}=\dfrac{1}{132}"

Construct the probability distribution of the random variable

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 & 3 & 4 & 5 \\ \hline
 p(x) & 1/132 & 5/44 & 25/66 & 25/66 & 5/44 & 1/132
\end{array}"